OpenStudy (anonymous):

HELP ASAP!!!!!!!!!!!!!!!!!!!!! A cylinder is inscribed in a cube. If the edge of the cube is 3" long, find the volume of the cylinder.

OpenStudy (anonymous):

 Radius of cylinder r=3/2 " Height h=3" \[Volume V=\pi r ^{2}h\] calculate V